OpenStudy (anonymous):

The ratio of blue marbles to yellow marbles in a jar is 7 to 9. If there are 28 blue marbles in the jar, how many yellow marbles are there? PLEASE, GIVE ME A REAL ANSWER TO THIS QUESTION.

OpenStudy (bibby):

There are 7 blue marbles for every 9 yellow The ratio is being multiplied by 4.

OpenStudy (anonymous):

can i have the answer so i can check my work?

OpenStudy (bibby):

Tell me your answer and I'll check it with /my/ work

OpenStudy (anonymous):

i got 36.

OpenStudy (bibby):

Yeah. Same. High five.

OpenStudy (anonymous):

are you sure its the right answer? :)

OpenStudy (bibby):

I totally graduated high school. I'll bet my life savings on this.

OpenStudy (anonymous):

it is....7/9 = 28/x (9 x 28)/7 = 252/7 = 36

OpenStudy (anonymous):

thank you soooo much :)

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

It's 36. because to find out how many times you need to times 7 to find 28 you divide it. 28 dived by 7 = 4. so it multiplies by four. so then you multiply 9 by 4 now. and you come up with 36!:)